OpenStudy (anonymous):

Solve x^2 + 27 = 0 and write factors. I got (x+9)(x+3) Is this one right?!?

OpenStudy (jdoe0001):

is \(\bf x^2\ or\ x^3?\)

OpenStudy (anonymous):

squared

OpenStudy (marissalovescats):

Aren't you missing your middle term?

OpenStudy (marissalovescats):

Is it solve x^2+12x+27?

OpenStudy (jdoe0001):

\(\bf x^2 + 27 = 0 \implies x^2=-27\implies x=\pm\sqrt{-27} \\ \quad \\ x=\pm\sqrt{27\cdot -1}\implies x=\pm\sqrt{27}\cdot \sqrt{-1} \\ \quad \\ x=\pm \sqrt{3^2\cdot 3}\cdot \sqrt{-1}\implies x=\pm 3\sqrt{3}\ i\)

OpenStudy (marissalovescats):

If not then ^^ that would make sense

OpenStudy (anonymous):

No.. That's what on the paper.

OpenStudy (jdoe0001):

well.... given +27.... it'd yield a negative -27 on the equation... maybe if you post a screenshot of the material

OpenStudy (anonymous):

I'm scanning the sheet now.

OpenStudy (anonymous):

It's upside down... Sorry

OpenStudy (anonymous):